\begin{thebibliography}{99}

\bibitem{Bernoulli}
%{Bernouilli}
%Якоб Бернулли, О законе больших чисел, М., Наука, 1986. 
%https://www.livelib.ru/book/1001201380-o-zakone-bolshih-chisel-yakob-bernulli
J. Bernoulli, 
Ars Conjectandi,
Basileae, Impensis Thurnisiorum, Fratrum, 1713. 
On the law of large numbers. Eng. transl. by O.B. Sheynin of the Pars Quarta of Ars Conjectandi, 2005; available at www.sheynin.de/download/bernoulli.pdf

\bibitem{Chebyshev}
%П.Л.Чебышевъ, О среднихъ величинахъ, Матем. сб., 2:1 (1867), 1–9.
P.L. Chebyshev [Tchebyshev], Des valuers moyennes, Journal de Math\'ematics Pures et Appliqu\'ees, 1867, vol. 12,  177-184.
%{\color{red}(указать в тексте теорему или формулу!)}

\bibitem{Borel1909}
%E. Borel, 
Émile Borel, Les probabilit\'es d\'enombrables et leurs applications arithm\'etiques. Rend. Circ. Matem. Palermo,  1909, 27, 247–271. %\\{\color{red}(chapitre II, section 11.)}
https://doi.org/10.1007/BF03019651

\bibitem{Cantelli}
F.P. Cantelli, Sulla probabilit\`a come limite della frequenza, Atti Accad. Naz. Lincei, 26(1) (1917), 39-45. Zbl 46.0779.02. Eng. transl.  
%by  C.F. Trustam and C.E. Bonferroni. 
in: 
C.F. Trustam and C.E. Bonferroni, Some recent Italian work on the Mathematical Basis of Actuarial Science, 
Journal of the Institute of Actuaries (1886-1994), 59(1), 
51-66. 
Published by: CUP on behalf of the Institute and Faculty of
Actuaries. 
On Probabilities considered as Limits of Frequencies. 
(A free translation of a Paper by F.P. Cantelli. Rome: R. Accademia dei Lincei, 1917.).  
57-61.
https://www.jstor.org/stable/41137221

%Also in [F.P. Cantelli, Alcune memorie matematiche. Onoranze Francesco Paolo Cantelli. Milan: Dott. A. Giuffrè, 222-229.???]

%{\color{red}(Lemma 2 \& formulae (28)-(29).)}

\bibitem{Mazurkiewicz}
S. Mazurkiewicz, O pewnem nog\'olneniu prawa wielkich liczb, 
Wiadomo\'sci matematycznych\footnote{ 
https://polona.pl/search/?query=stefan\_mazurkiewicz\&filters=public:1}, 1917, XXII, 69-82.



%\bibitem{Petrov}-?
%Valentin V. Petrov, On the strong law of large numbers, Statistics \& Probability Letters, 1966,  26, 377-380.



%\bibitem{Schach}
%I. Karatzas, W. Schachermayer, A Strong Law of Large Numbers for Positive Random Variables,arXiv:2111.15469 https://doi.org/10.48550/arXiv.2111.15469

\bibitem{Rajchman32}
A. Rajchman, Zaostrzone prawo wielkich liczb, Mathesis Polska\footnote{http://mbc.cyfrowemazowsze.pl/dlibra/applet?mimetype=image/x.djvu\&sec=false\\\&handler=djvu\_html5\&content\_url=/Content/15684/index.djvu}, 
%6, 
1931, VI 
%Rok. 6, 
(7-8), 145 - 161.
%(the case with finite variances but for pairwise independent r.v.)
%(see Chandra3 -- 91.)
%http://mbc.cyfrowemazowsze.pl/dlibra/applet?mimetype=image/x.djvu\&sec=false\&handler=djvu\_html5\&content\_url=/Content/15684/index.djvu

\bibitem{Chandra3}
T.K. Chandra, Extensions of Rajchman’s Strong Law of Large Numbers. Sankhyā: The Indian Journal of Statistics, Series A (1961-2002), 1991, 53(1), 118–121. http://www.jstor.org/stable/25050825

\bibitem{Kolmogorov}
A.N.Kolmogorov, 
Foundations of the Theory of Probability, 2nd Russian edition, Moscow, Nauka, 1974 (in Russian).
%NB: the author was unable to find any English translation from the second Russian edition; all translations into English found were made well before 1974.
%А.Н.Колмогоров, 
%Основные понятия теории вероятностей, М., Главная редакция физ.-матем. литературы изд-ва ``Наука'', изд. 2, 1974. 

\bibitem{Etemadi81}
N. Etemadi, An elementary proof of the strong law of large numbers, Z. Wahrscheinlichkeitstheorie und verw. Gebiete, 1981, 55, 119 - 122.

\bibitem{Hungarians}
S. Cs\"orgo, K. Tandori, V. Totik, 
On the Strong Law of Large Numbers for Pairwise Independent Random Variables, Acta Mathematica Hungarica, Vol. 42, No. 3 - 4, 1983, pp. 319-330. doi:10.1007/BF01956779 

\bibitem{Rogge}
D. Landers, L. Rogge,  Laws of large numbers for pairwise independent uniformly
integrable random variables. Math. Nachr., 1987, 130, 189 - 192.

\bibitem{Rogge2}
D. Landers, L. Rogge, 
Laws of large numbers for uncorrelated Ces\`aro uniformly integrable random variables, 
Sankhy$\bar{\text{a}}$: The Indian Journal of Statistics, 
1997, 59, Series A, Pt. 3,  301 - 310. 
%$\bar a$

\bibitem{BoseChandra94}
A. Bose, T.K. Chandra, A Note on the Strong Law of Large Numbers. Calcutta Statistical Association Bulletin,  1994, 44(1–2), 115 - 122. https://doi.org/10.1177/0008068319940109






\bibitem{Chandra}
T.K. Chandra, 
Uniform Integrability in the Cesàro Sense and the Weak Law of Large Numbers, 
Sankhyā: The Indian Journal of Statistics, Series A (1961-2002), 1989, 
51(3), 309 - 317.  
https://www.jstor.org/stable/25050754


\bibitem{Chandra2}
T.K. Chandra, A. Goswami, 
Ces\'aro Uniform Integrability and the Strong Law of Large Numbers, 
Sankhyā: The Indian Journal of Statistics, Series A (1961-2002), 1992, 
54, 215 - 231. 
https://www.jstor.org/stable/25050875

\bibitem{Chandra_book}
T.K. Chandra, Laws of large numbers, Narosa, New Delhi et al., 2012.

\bibitem{Seneta92}
E. Seneta, 
On the History of the Strong Law of Large Numbers
and Boole’s Inequality, 
Historia Mathematica, 
1992, 19(1), 24 - 39.

\bibitem{Seneta3}
 E. Seneta. A Tricentenary history of the Law of Large Numbers. Bernoulli, 2013, 19 (4), 1088 - 1121. https://doi.org/10.3150/12-BEJSP12 

\bibitem{Shiryaev}
%А.Н. Ширяев, К 200-летию со дня рождения великого русского математика П.Л. Чебышёва, Теория вероятн. и ее примен., 2021, 66(4), 625–635. https://doi.org/10.4213/tvp5523
A.N. Shiryaev, On the bicentenary of the birth of P. L. Chebyshev, a great Russian mathematician”, Theory Probab. Appl., 66:4 (2022), 497 - 505
https://doi.org/10.1137/S0040585X97T990575


\bibitem{AA_AV24}
 A. T. Akhmiarova and A. Yu. Veretennikov, On Strong Law of Large Numbers for Pairwise Independent Random Variables, Theory of Probability and its Applications, 2024, Volume 69, Issue 3, Pages 337 - 345 DOI: https://doi.org/10.1137/S0040585X97T991969
%	arXiv:2211.04435, 
%https://doi.org/10.48550/arXiv.2211.04435



%\bibitem{bingham86}
%N.H. Bingham, Extensions of the Strong Law, Advances in Applied Probability, 1986, 18, 27–36. http://www.jstor.org/stable/20528774


\bibitem{MZ}
J. Marcinkiewicz, A. Zygmund, Sur les fonctions ind\'ependantes, Fund. Math., 1937, 28, 60-90. 
Reprinted in Józef Marcinkiewicz, Collected papers, edited by Antoni Zygmund, Panstwowe Wydawnictwo Naukowe, Warsaw, 1964, 233 - 259.

\bibitem{Sawyer69}
S. Sawyer, 
Maximal Inequalities of Weak Type, 
Annals of Mathematics, Second Series, 1966, 84(2), 
157 - 174. 
%Published by: Mathematics Department, Princeton University
https://www.jstor.org/stable/1970516




%\bibitem{Janisch}-?
%M. Janisch, Kolmogorov's strong law of large numbers holds for pairwise uncorrelated random variables, 
%Теория вероятн. и ее примен., 66:2 (2021), 327–341; 
%Theory Probab. Appl., 66:2 (2021), 263–275. https://doi.org/10.4213/tvp5459

%\bibitem{Khintchin}-?
%A.Ya. Khinchine, Sur la loi des grands nombres, Comptes Rendus Acad. Sci. Paris, 1929, vol.1929/01 (T188) - 1929/06, 477-479.

%\bibitem{Khintchin2}A.Ya.Khinchine 36? 



\bibitem{Korchevsky}
%В.М. Корчевский, Об усиленном законе больших чисел для последовательностей попарно независимых случайных величин, Вероятность и статистика. 26, Зап. научн. сем. ПОМИ, 466, ПОМИ, СПб., 2017, 159–166.
V.M. Korchevsky, On the Strong Law of Large Numbers for sequences of pairwise independent random variables, J. Math. Sci., 2020, 244, 805 - 810. https://doi.org/10.1007/s10958-020-04654-y



\bibitem{Shiryaev2}
A.N. Shiryaev, Probability (2nd ed.). Springer, 1996. 
%(Lemma IV.3.2)

%\bibitem{ZSCh}-
%A.M. Zubkov, B.A. Sevastyanov, V.P. Chistiakov, Exercises on probability theory, Moscow, Lan', 2021 (in Russian)
%А.М. Зубков, Б.А. Севастьянов, В.П. Чистяков, Сборник задач по теории вероятностей, М., Лань, 2021. 
%Задача 4.24

\end{thebibliography}

\documentclass[12pt,a4paper]{article}

\usepackage[T2A]{fontenc}
\usepackage{mathrsfs}

\usepackage[utf8]{inputenc}
\usepackage[english]{babel}
\usepackage{indentfirst}
\usepackage{misccorr}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{hyperref}
\usepackage{color}

\newif\ifru
\rufalse

\newif\ifen
\entrue % \*true or \*false
\newtheorem{theorem}{Theorem}

\newtheorem{Lemma}{Lemma}

\ifen
\title{{\normalsize\tt\hfill\jobname.tex}\\
On strong law of large numbers for non identically distributed random variables
}

\author{I.V. Kozlov\footnote{M.V. Lomonosov Moscow State University; Institute for Information Transmission Problems (A.A.Kharkevich Institute); Vega Institute Foundation; email: ilia.kozlov@math.msu.ru}, \; A.Yu. Veretennikov\footnote{Institute for Information Transmission Problems (A.A.Kharkevich Institute); RUDN University; NRU HSE; MSU; email: alexander.veretennikov2011@ya.ru}
}
\fi

\ifru
\title{{\normalsize\tt\hfill\jobname.tex}\\
Об усиленном  законе больших чисел для попарно независимых случайных величин
}

\author{И.В. Козлов\footnote{Московский государственный университет им. М.В. Ломоносова; Институт проблем передачи информации им. А.А. Харкевича; email: } \; А.Ю. Веретенников\footnote{Институт проблем передачи информации им. А.А. Харкевича; РУДН ; email: alexander.veretennikov2011@ya.ru}
}
\fi

\date{}

\begin{document}
\maketitle

\ifen
\begin{abstract}
\noindent
%Some new remarks concerning strong law of large numbers (SLLN) for pairwise independent random variables are presented. The main point is relaxing the assumption on the existence of a mean value for each random variable.

A new  version of a strong law of large numbers for a ``good'' pairwise independent sequence of random variables (r.v.'s) with a small part of ``bad'' dependent r.v.'s is proposed. The main goal is to relax the assumption on the existence of the expectation for each summand: the members of an ``infrequent'' part of the whole sequence may have moments of orders converging to zero.

\medskip

\noindent
Keywords: Strong law of large numbers; partial pairwise independence; partal dependence; partial absence of expectations.

\medskip

\noindent
MSC2020: 60F15

\end{abstract}
\fi

\ifru
\begin{abstract}
\noindent

Предложена новая версия усиленного закона больших чисел для ``хорошей'' последовательности случайных величин (с.в.) с небольшой частью ``плохих'' зависимых с.в. Основной целью являлось ослабление требования на существование математического ожидания для каждого слагаемого: члены ``разреженной'' части всей последовательности может иметь моменты порядков, сходящихся к нулю.

\medskip

\noindent
Ключевые слова: усиленный закон ольших чисел; попарная независимость; ччастичная зависимость; частичное отсутствие математических ожиданий.

\medskip

\noindent
MSC2020: 60F15

\end{abstract}
\fi

\ifen
\section{Introduction}
The Law of Large Numbers (LLN for what follows) is the first rigorously proved limit theorem in probability. The interest to any new its version is natural because the majority of  methods of parameter estimations and hypothesis testing is based on LLN, but also due to a philosophical importance of such theorems. The history of the LLN goes back to the work by Jakob Bernoulli in 1713, see \cite{Bernoulli},  then continued by de Moivre, Laplace, Chebyshev \cite{Chebyshev} et al. In the 20th century the strong form of an LLN was discovered by Borel \cite[chapter II, section 11]{Borel1909}, Cantelli  \cite[Lemma 2 \& formulae (28)-(29)]{Cantelli}, Mazurkiewicz  \cite{Mazurkiewicz}, Rajchman \cite{Rajchman32} (see also a recent exposition of the latter work in \cite{Chandra3}), which was brought to a nearly final form by Kolmogorov \cite{Kolmogorov}, yet further developed by Etemadi  \cite{Etemadi81}, Cs\"orgo, Tandori, and Totik \cite{Hungarians}, Landers and Rogge \cite{Rogge, Rogge2}, Chandra \cite{BoseChandra94, Chandra, Chandra2, Chandra_book, Chandra3}, among others. The history may be recalled in \cite{Seneta92, Seneta3, Shiryaev}.

The goal of the present note is to offer some further  extension for the case of pairwise independent random variables, which includes Kolmogorov's SLLN \cite[theorem 6.5.3]{Kolmogorov}, as well as the recent result in \cite{AA_AV24}. The matter is that a certain, not necessarily finite part of the summands may have infinite expectations with moments of orders, which may converge to zero. The most important technical steps relate to the calculi in \cite{Kolmogorov, MZ, Sawyer69}. 

The present paper consists of four sections: this introduction, the setting, the SLLN for partially pairwise independent random variables (theorem 
\ref{thm1}), the proof of theorem \ref{thm1}.
\fi

\ifen
\section{The setting} %Классические результаты}
Let us consider a sequence $(X_n, n\ge 1)$ of pairwise independent random variables (r.v's) with finite expectations, for which without loss of generality the equalities 
\begin{equation}\label{eq2-1}
\mathsf E X_i = 0, \qquad i\ge 1,
\end{equation}
hold. It is also assumed that there exists $n_0 \ge 1$ such that 
\begin{equation}\label{CUI}
C(n_{0})
  := \int_{0}^{\infty}
       \sup_{n \ge n_{0}}
       \frac{1}{n}
       \sum_{k = n_{0}}^{n}
      \mathsf P\!\bigl(|X_{k}| > x\bigr)\,dx
    < \infty.
\end{equation}

Moreover, there is another sequence of r.v's $(Y_n)$, arbitrarily dependent and, in general, without finite expectations. Instead, it is assumed that there is a nonrandom sequence of values $(a_n)$, such that all $a_n>0$ and $a_n\downarrow 0, \, n\to\infty$, and $\mathsf E |Y_n|^{a_n}<\infty$; an additional technical condition on the sequence $(Y_n)$ is as follows: let 
\begin{equation}\label{defVn}
    V_n := |Y_n|^{a_n}, \quad \mathsf EV_n < \infty,
\end{equation}
and there exists a nonincreasing function $\bar G(t), \, t\ge0$ such that 
\begin{equation}\label{VnCG}
  \sup_{n\ge 1}\mathsf P\bigl(V_n>t\bigr)\le\bar G(t), \qquad
  \int_{0}^{\infty}\bar G(t)dt =: C_G <\infty.
\end{equation}

Now, the sequence of r.v's $(Z_n)$ is constructed as a mixture of $(X_n)$ and $(Y_n)$. More precisely, there is one more nonrandom sequence $(\alpha_n)$ taking values $0$ and $1$; denote $\varphi_n:= \sum_{k=1}^n \alpha_k$ and $\psi_n:= n - \varphi_n$, and, finally, define 
$$
Z_n:= Y_{\varphi_n} 1(\alpha_n = 1) + X_{\psi_n} 1(\alpha_n = 0).
$$
The law of large numbers (LLN in what follows) will be stated and proved for the sequence $(Z_n)$ under the standing condition that the sequence $(Y_n)$ is ``infrequent'', see what follows. Notice that recently a strong LLN was established in a likewise case under a more restrictive condition that all $a_n\equiv a \in (0,1)$, see \cite{AA_AV24}.

The condition of ``infrequency'' of the sequence $(Y_n)$ is as follows:

\begin{equation}\label{finan}
\limsup_{n \to \infty} \frac{\varphi_n^{\frac{1}{a_{\varphi_n}}}}{n} < \infty \quad (\text{i.e., } \exists \; C>0: \; \varphi_n^{\frac{1}{a_{\varphi_n}}} \le Cn \; \text{for  $n$ large enough.}) \end{equation}
Naturally, we wish that the sequence $(\varphi(n))$ increase to infinity (otherwise it is not an interesting case):
\begin{equation}\label{anlnn}
\varphi(n) \to \infty, \quad n\to\infty. 
\end{equation}
The condition (\ref{anlnn}) will also be assumed, although, formally, the result will hold for a bounded sequence $(\varphi(n))$, too.

\begin{theorem}\label{thm1}
The assumptions (\ref{eq2-1})–(\ref{anlnn}) imply an SLLN 
\begin{equation}\label{thmSLLN}
\frac{1}{n}\sum_{k=1}^{n} Z_k \xrightarrow{\text{a.s.}} 0.
\end{equation}
\end{theorem}
\fi

\ifen
\section{Proof of Theorem}
We will need two lemmata. One is a simple analogue of N. Etemadi's theorem from  \cite{Etemadi81} under relaxed Cesàro UI condition, the other is an advanced analogue of the a calculus  from \cite{Sawyer69}, see also \cite{Korchevsky}.

\begin{Lemma}[for $(X_n)$]\label{lem:X}
Let the assumptions (\ref{eq2-1})–-(\ref{CUI}) be satisfied. Then
\begin{equation}\label{le1-eq1}
\frac{1}{n}\sum_{k=1}^{n-\varphi(n)} X_k \xrightarrow{\text{a.s.}} 0,
\qquad n\to\infty.
\end{equation}
\end{Lemma}

\begin{proof}
First of all, we have for each $n_0$, 
\[
\frac{1}{n}\sum_{k=1}^{n_0-1} X_k \xrightarrow{\text{a.s.}} 0.
\qquad n\to\infty,
\]
Hence, it suffices to verify \[
\frac{1}{n}\sum_{k=n_0}^{n-\varphi(n)} X_k \xrightarrow{\text{a.s.}} 0,
\qquad n\to\infty.
\]
%{\color{red}The latter follows from theorem 1 in \cite{Etemadi81}.}
The latter follows from the results of \cite{BoseChandra94}; see also Theorem~2 in \cite{Korchevsky}.
\end{proof}

\begin{Lemma}[for $(Y_n)$]\label{lem:Y}
Let assumptions  (\ref{defVn}) -- (\ref{finan})  be satisfied. Then
\begin{equation}\label{le2-eq1}
\frac{1}{n}\sum_{k=1}^{\,\varphi(n)}Y_k \xrightarrow{\text{a.s.}} 0,
\qquad n\to\infty.
\end{equation}
\end{Lemma}
\begin{proof} Let $\tilde V_n := \min\{V_n, n\}$. Let us show that for any $p>1$, 
\[
  \mathsf E\sum_{n=3}^{\infty}\frac{\tilde V_n^{p}}{n^{p}} < \infty.
\]
We have, 
\begin{align*}
& \mathsf E\tilde V_n^{p}
  =\int_{0}^{n^{p}}
    \mathsf  P\bigl(\tilde V_n^p>u\bigr)du
  =\int_{0}^{n} p s^{p-1} \mathsf P(V_n>s)ds
   + n^{p}  \mathsf P(V_n>n)
   \\\\
&  \le
  \int_{0}^{n} p s^{p-1}\bar G(s)ds
  + n^{p}\bar G(n).
\end{align*}
Therefore, 
\[
  \sum_{n=3}^{\infty}\frac{\mathsf E\tilde V_n^p}{n^{p}}
  \le
  A+B,
\]
where it was denoted
\[
  A:=\sum_{n=3}^{\infty}\frac1{n^{p}}
      \int_{0}^{n}p s^{p-1}\bar G(s)ds,
  \qquad
  B:=\sum_{n=3}^{\infty}\bar G(n).
\]

Let us evaluate for $p>1$ the remainder of the series $\sum_{n=m}^{\infty}n^{-p}$:
\[
\sum_{n=m}^{\infty} \frac{1}{n^p} \le \int_{m-1}^{\infty} \frac{1}{x^p}\,dx = \frac{(m-1)^{1-p}}{p - 1}, \qquad p > 1,\ m \ge 2.
\]
Hence, 
\begin{align*}
&  A =\sum_{n=3}^{\infty}\frac1{n^{p}}\sum_{i=1}^{n} \int^{i}_{i-1}s^{p-1} p \bar G(s) ds = \sum_{i=1}^{\infty} \left( \int_{i-1}^{i} s^{p-1} p \, \bar{G}(s) \, ds \right) \sum_{n=\max(i,3)}^{\infty} \frac{1}{n^p}
 \\\\
&  \le \sum_{i=1}^{\infty} \left( \int_{i-1}^{i} i^{p-1} p \, \bar{G}(s) \, ds \right) \sum_{n=\max(i,3)}^{\infty} \frac{1}{n^p} = \Bigg[\sum_{i=3}^{\infty} \left( \int_{i-1}^{i} i^{p-1} p \, \bar{G}(s) \, ds \right) \sum_{n=i}^{\infty} \frac{1}{n^p} 
 \\\\
& +  \left( \int_{1}^{2} 2^{p-1} p \, \bar{G}(s) \, ds \right) \sum_{n=3}^{\infty} \frac{1}{n^p} + \left( \int_{0}^{1} p \, \bar{G}(s) \, ds \right) \sum_{n=3}^{\infty} \frac{1}{n^p}\Bigg]
 \\\\
&    \le \Bigg[\sum_{i=3}^{\infty} \left( \int_{i-1}^{i} i^{p-1} p \, \bar{G}(s) \, ds \right) \frac{(i-1)^{1-p}}{p - 1} + \left( \int_{1}^{2} 2^{p-1} p \, \bar{G}(s) \, ds \right) \frac{(2)^{1-p}}{p - 1}
 \\\\
& + \left( \int_{0}^{1} p \, \bar{G}(s) \, ds \right) \frac{(2)^{1-p}}{p - 1}\Bigg]
 %\\\\& 
\le \frac{p}{p-1}\int^{\infty}_{0} \bar G(s)ds \le \frac{p}{p-1} C_G.
\end{align*}
Since the function $\bar G$ is non increasing, we get 
\(\displaystyle  \bar G(n)\le\int_{n-1}^{n}\bar G(t)dt\). 
So, 
\[
B\le\int_{2}^{\infty}\bar G(t)dt< C_G.
\]
Thus, we obtain, 
\[
  \mathsf E\sum_{n=3}^{\infty}\frac{\tilde V_n^{p}}{n^{p}}\le C_G\Bigg(\frac{p}{p-1}+1\Bigg) = C_G \Bigg(\frac{2p-1}{p-1}\Bigg),
  \qquad p>1.
\]

Let $p_k:=1/a_k>1$. Then
\[
 \mathsf  E\sum_{n=3}^{\infty}\frac{\tilde V_n^{p_k}}{n^{p_k}} \le C_G \Bigg(\frac{2p_k-1}{p_k-1}\Bigg),
  \quad\forall k\in\mathbb N.
\]

Let us choose a sequece of numbers $\{N_k\}_{k \in \mathbb N}$ so as $N_{k+1} \ge N_k+1$ and
\[
 \mathsf E\sum_{n=N_k+1}^{\infty}\frac{\tilde V_n^{p_k}}{n^{p_k}}
  <\frac1{k^{2}}.
\]
Such a sequence does exist because each series converges and its remainder goes to zero. Let 
\[
  q(n):=p_k,\quad \text{if} \quad N_k < n\le N_{k+1}.
\]
Then we have, 
\[
 \mathsf E\sum_{n=3}^{\infty}\frac{\tilde V_n^{q(n)}}{n^{q(n)}}
  \le
  \underbrace{ \mathsf E\sum_{n=3}^{N_1}\frac{\tilde V_n^{p_1}}{n^{p_1}}}_{C(p_1)}
  +\sum_{k=1}^{\infty}\frac1{k^{2}}
  <C+\frac{\pi^{2}}{6}<\infty.
\]
So, 
\[
   \mathsf  E\sum_{n=3}^{\infty}\frac{\tilde V_n^{q(n)}}{n^{q(n)}} < \infty
\]
Since $\frac{\tilde V_n}{n} \le 1$ and $q(n)\le p_n$, it follows that  $\displaystyle \Big[\frac{\tilde V_n}{n}\Big]^{q(n)} \ge \Big[\frac{\tilde V_n}{n}\Big]^{p_n}$. Therefore,
\[
 \mathsf E\sum_{n=3}^{\infty}\frac{\tilde V_n^{p_n}}{n^{p_n}} \le  \mathsf  E\sum_{n=3}^{\infty}\frac{\tilde V_n^{q(n)}}{n^{q(n)}} < \infty
\]
This implies convergence a.s. 
\[
  \sum_{n=3}^{\infty}\frac{\tilde V_n^{p_n}}{n^{p_n}}<\infty
  \quad(\text{a.s.}).
\]
Recall that $\tilde V_n=\min(V_n,n)$. Therefore, we get, 
\[
  \sum_{n=1}^{\infty}\frac{V_n^{p_n}}{n^{p_n}}
  =\sum_{n=1}^{\infty}\frac{|Y_n|^{a_n p_n}}{n^{p_n}}<\infty
  \quad(\text{a.s.}).
\]
Now, let $p_n=1/a_n$. Then we obtain 
\[
  \sum_{n=1}^{\infty}\frac{|Y_n|^{a_n/a_n}}{n^{1/a_n}} = \sum_{n=1}^{\infty}\frac{|Y_n|}{n^{1/a_n}}<\infty
  \quad(\text{a.s.}).
\]
By virtue of the Kronecker lemma (see, for example, \cite[Lemma IV.3.2]{Shiryaev2}), 
\[
\frac{1}{n^{\frac1{a_n}}} \sum_{k=1}^{n} |Y_k| \xrightarrow{\text{a.s.}} 0, \quad n\xrightarrow{}\infty
\]
Let $n = \varphi(m)$, then by virtue of the conditions (\ref{finan}) and (\ref{anlnn}) we get, 
\[
\frac{1}{m}
\big|\sum_{k=1}^{\varphi(m)} Y_k\big| \le  
\frac{C}{\varphi(m)^{1/a_{\varphi(m)}}}
\sum_{k=1}^{\varphi(m)} |Y_k| \xrightarrow{\text{a.s.}} 0, \quad m\xrightarrow{}\infty.
\]
The lemma is proved. 
\end{proof}
\begin{proof}[Proof of theorem]
The goal is to show convergence
\[
  \mathsf{P}\Bigl(\frac{1}{n}\sum_{k=1}^n Z_k \to 0,\; n\to\infty\Bigr) = 1.
\]
It follows from (\ref{le1-eq1}) and (\ref{le2-eq1}). The theorem is proved.
\end{proof}
\fi

\ifen
\section*{Acknowledgements}
For the second author this work was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics ``BASIS''.

\fi

\ifru
\section*{Благодарности}
For the second author this work was supported by the Foundation for the advancement of theoretical physics and mathematics ``BASIS''.

\fi

\ifen

\fi

\ifru

\fi

\end{document}